A sum of 6000 is invested, part of it at 9% interest and the remainder at 11%. If the interest earned by the 9% investment is 160 less than the interest earned by the 11% investment, find the amount invested at each rate

Answer by addingup(3677)

0.09x = 0.11(6000-x)-160
0.09x = 660-0.11x-160
0.20x = 500
x = 2500 this is the amount invested at 9%
6000-2500 = 3500 is the amount invested at 11%
-----------------------
check:
0.11*3500 = 385
0.09*2500 = 225
. . . . . -----
Difference: 160 Correct
:
John